### v1

# The weakest condition for a Kannan-type contraction mapping on a complete metric space

Abstract.
In this paper, we investigate the weakest possible conditions for fixed point theorems concerning two types of mappings: Kannan and Chatterjea. We employ the so-called CJM condition, which has been successfully applied to the classical Banach-type mappings by Ćirić [7]. We show that the CJM condition is nearly the weakest possible to have a fixed point, and that a slight modification yields the weakest condition even for these two types of mappings.
Key words and phrases:
fixed point theorem, Kannan-type contraction, Chatterjea-type contraction
2020 Mathematics Subject Classification:
54H25, 47H10
1. Introduction
Fixed-point theory is an important area in nonlinear analysis, providing foundational results for solving various mathematical and applied problems. Among the classical contractive conditions, Kannan-type mappings have been widely studied due to their unique properties and applications. In 1968 Kannan [10] introduced a contraction condition that does not require strict continuity like Banach [2] contractions. A mapping $T:X\to X$ in a metric space $(X,d)$ is called a Kannan contraction if there exists a constant $\alpha\in\left[0,\frac{1}{2}\right)$ such that:
$d(Tx,Ty)\leq\alpha\left(d(x,Tx)+d(y,Ty)\right),\text{for all}\ x,y\in X.$
Several authors [11, 12, 15, 4] have studied generalizations of Kannan mappings, including integral-type conditions and applications in various spaces. In 1974, L. B. Ćirić [7] introduced the generalization of Banach’s contraction principle, which laid the groundwork for what would later become known as CJM contractions. Building on this foundation, J. Matkowski [13] contributed significantly to the development of the CJM framework in 1980 by studying fixed point theorems for contractive mappings in metric spaces, solidifying the theoretical basis for this class of contractions. Finally, J. Jachymski [9] provided a comprehensive study of contractive conditions in 1995, introducing the concept of equivalent conditions for Meir-Keeler-type theorems [14] and further advancing the understanding of generalized contractions.
CJM contractions provide a powerful and unified framework for analyzing fixed point theorems in metric spaces. Specifically, a mapping $T$ on a metric space $(X,d)$ satisfies the CJM condition if the following holds for all $x,y\in X$:
  1. (i)
    For every $\varepsilon>0$, there exists $\delta>0$ such that $d(x,y)<\varepsilon+\delta$ implies $d(Tx,Ty)\leq\varepsilon$.
  2. (ii)
    $x\not=y$ implies $d(Tx,Ty)<d(x,y)$.
T. Suzuki [17] introduced the Suzuki-type contractions in 2008 which is a generalization that weakens the Banach contraction condition. It not only guarantees fixed points but also characterizes the completeness of the underlying metric space. In 2017, T. Suzuki [18] established significant results using CJM contractions and provided a unified framework for characterizing various contractive conditions, such as Meir-Keeler, Boyd-Wong, and Browder contractions, by using the convergent sequences. However, the article does not explicitly focus on Kannan-type mappings to prove results within the CJM framework. These studies collectively advance our understanding of Kannan-type mappings by developing new analytical techniques and expanding the theoretical scope through generalized contractive conditions see for instance [1, 16, 8].
In this article, Kannan-type mapping has been explored using CJM contractions to generalize the fixed point results. It is achieved using the weakest condition of T.Suzuki’s [18] work to prove fixed point results for Kannan-type mappings as previously not covered by this approach. We also showed the same for Chatterjea-type contraction mapping [6], which is similar to Kannan-type mapping.
2. The CJM version of a Kannan-type contraction mapping
The following fixed point theorem is proved in preparation for stating the weakest condition.
Theorem 2.1.
Let $(X,d)$ be a complete metric space and suppose that the mapping $T:X\to X$ satisfies the following.
  1. (i)
    For every $\varepsilon>0$, there exists $\delta>0$ such that $\frac{1}{2}\{d(x,Tx)+d(y,Ty)\}<\varepsilon+\delta$ implies $d(Tx,Ty)\leq\varepsilon$.
  2. (ii)
    $x\not=y$ implies $d(Tx,Ty)<\frac{1}{2}\{d(x,Tx)+d(y,Ty)\}$.
Then, $T$ has a unique fixed point.
Proof.
Fix $x_{0}\in X$. Set $x_{n}=T^{n}x_{0}$ for all $n\in\mathbb{N}$. If there exists $N\in\mathbb{N}$ such that $x_{N}=x_{N+1}$, $x_{N}$ is a fixed point of $T$. Therefore, assume that $x_{n}\not=x_{n+1}$ for any $n\in\mathbb{N}$. Then, we have
$d(x_{n},x_{n+1})$ $=d(Tx_{n-1},Tx_{n})<\frac{1}{2}\{d(x_{n-1},Tx_{n-1})+d(x_{n},Tx_{n})\}$
$=\frac{1}{2}\{d(x_{n-1},x_{n})+d(x_{n},x_{n+1})\}.$
So,
$\frac{1}{2}d(x_{n},x_{n+1})<\frac{1}{2}d(x_{n-1},x_{n}).$
Therefore, $\{d(x_{n},x_{n+1})\}$ is a strictly decreasing sequence.
Thus, there exists $\varepsilon_{0}\geq 0$ such that $d(x_{n},x_{n+1})\to\varepsilon_{0}$ as $n\to\infty$ and $d(x_{n},x_{n+1})>\varepsilon_{0}$ for any $n$.
Assume that $\varepsilon_{0}>0$. Then there exist $\delta_{0}>0$ and $k\in\mathbb{N}$ such that
$\frac{1}{2}\{d(x_{k},Tx_{k})+d(x_{k+1},Tx_{k+1})\}<\varepsilon_{0}+\delta_{0}.$
By assumption (i) we have $d(Tx_{k},Tx_{k+1})\leq\varepsilon_{0}$, which is a contradiction. Therefore, $\varepsilon_{0}=0$ and
(2.1) $\lim_{n\to\infty}d(x_{n},x_{n+1})=0.$
Next, we show that $\{x_{n}\}$ is a Cauchy sequence. Let $\varepsilon>0$ be arbitrary and let $\delta$ be chosen such that $\delta\leq\varepsilon$. By (2.1) there exists some $k\in\mathbb{N}$ such that $d(x_{n-1},x_{n})<\delta$ for each $n>k$. Fix $n>k$. It suffices to show that
$d(x_{n},x_{n+p})\leq\varepsilon,\ \text{for all}\ p\in\mathbb{N}.$
From $d(x_{n},x_{n+1})\leq\delta$ by (2.1), we have
$\frac{1}{2}\{d(x_{n-1},Tx_{n-1})+d(x_{n+p-1},Tx_{n+p-1})\}$ $=\frac{1}{2}\{d(x_{n-1},x_{n})+d(x_{n+p-1},x_{n+p})\}$
$\leq\delta<\varepsilon+\delta.$
Then by assumption (i), we obtain
$d(x_{n},x_{n+p})=d(Tx_{n-1},Tx_{n+p-1})\leq\varepsilon.$
Therefore since $\{x_{n}\}$ is a Cauchy sequence, there exists $u\in X$ such that $x_{n}\to u$ as $n\to\infty$. From
$d(u,Tu)$ $\leq d(u,x_{n})+d(x_{n},Tx_{n})+d(Tx_{n},Tu)$
$<d(u,x_{n})+d(x_{n},Tx_{n})+\frac{1}{2}\{d(x_{n},Tx_{n})+d(u,Tu)\},$
we have
$\frac{1}{2}d(u,Tu)<d(u,x_{n})+\frac{3}{2}d(x_{n},Tx_{n})\to 0\ (n\to\infty).$
Then we get $u=Tu$.
Finally, we show uniqueness. Let $u\not=v,\ Tu=u$ and $Tv=v$. Then we have
$d(u,v)=d(Tu,Tv)<\frac{1}{2}\{d(u,Tu)+d(v,Tv)\}=0,$
which is a contradiction. This completes the proof of the theorem. ∎
3. The weakest condition for a Kannan-type contraction mapping
The weakest condition is a slightly weakened version of the condition for contraction mapping in Section 2.
Theorem 3.1.
Let $(X,d)$ be a complete metric space and suppose that the mapping $T:X\to X$ satisfies the following.
(CM) $x\not=y\ \text{implies}\ d(Tx,Ty)<\frac{1}{2}\{d(x,Tx)+d(y,Ty)\}.$
Then, the following are equivalences.
  1. (i)
    For any $x\in X$ and $\varepsilon>0$, there exists $\delta>0$ such that
    $\frac{1}{2}\{d(T^{i}x,T^{i+1}x)+d(T^{j}x,T^{j+1}x)\}<\varepsilon+\delta$ implies $d(T^{i+1}x,T^{j+1}x)\leq\varepsilon$ for all $i,j\in\mathbb{N}\cup\{0\}$.
  2. (ii)
    $T$ has a unique fixed point $z\in X$ and $\{T^{n}x\}$ converges to $z$ for any $x\in X$.
Proof.
((i)$\Rightarrow$(ii)) It can be shown as in the proof of Theorem 2.1.
((ii)$\Rightarrow$(i)) Let $z\in X$ be a fixed point of $T$. Arguing by contradiction, we assume that (i) does not hold, that is, there exist $x\in X$ and $\varepsilon>0$ such that for all $n\in\mathbb{N}$ there exists $\{f(n)\},\{g(n)\}\subset\mathbb{N}\cup\{0\}$ such that
$f(n)<g(n),$
(3.1) $\frac{1}{2}\{d(x_{f(n)},x_{f(n)+1})+d(x_{g(n)},x_{g(n)+1})\}<\varepsilon+\frac{1}{n}\quad\text{and}$
(3.2) $\varepsilon<d(x_{f(n)+1},x_{g(n)+1}),$
where $x_{f(n)}=T^{f(n)}x$. Without loss of generality, we may assume that either
  • $\lim_{n\to\infty}f(n)=\infty$  or
  • There exists $\mu\in\mathbb{N}\cup\{0\}$ such that $f(n)=\mu$ for all $n\in\mathbb{N}$.
Note that if $\{f(n)\}$ oscillates, the second case is attributed. Similarly, we may assume that either
  • $\lim_{n\to\infty}g(n)=\infty$  or
  • There exists $\nu\in\mathbb{N}\cup\{0\}$ such that $g(n)=\nu$ for all $n\in\mathbb{N}$.
By (3.2) we have $x_{f(n)}\not=x_{g(n)}$. So, by (CM) we obtain
$\varepsilon$ $<d(x_{f(n)+1},x_{g(n)+1})<\frac{1}{2}\{d(x_{f(n)},x_{f(n)+1})+d(x_{g(n)},x_{g(n)+1})\}<\varepsilon+\frac{1}{n}.$
Therefore, we get
(3.3) $0<\varepsilon=\lim_{n\to\infty}d(x_{f(n)+1},x_{g(n)+1})=\frac{1}{2}\lim_{n\to\infty}\{d(x_{f(n)},x_{f(n)+1})+d(x_{g(n)},x_{g(n)+1})\}.$
Since $\{x_{n}\}$ converges to $z$, $f(n)\not\to\infty$ or $g(n)\not\to\infty$ holds. Suppose that $f(n)\not\to\infty$ holds, that is, $f(n)=\mu$ for all $n\in\mathbb{N}$ is satisfied.
First, if $g(n)=\mu$ for all $n\in\mathbb{N}$, we have
$d(x_{\mu+1},x_{\nu+1})=\varepsilon,$
which is a contradiction. On the other hand, if $\lim_{n\to\infty}g(n)=\infty$, then $x_{g(n)}\to z$, so from (3.3) we have
(3.4) $d(x_{\mu+1},z)=\varepsilon$ $=\frac{1}{2}d(x_{\mu},x_{\mu+1})\leq\frac{1}{2}\{d(x_{\mu},z)+d(z,x_{\mu+1})\}.$
We can show the first equality in (3.4) since we obtain
$|d(x_{\mu+1},x_{g(n)+1})-d(x_{\mu+1},z)|\leq d(x_{g(n)+1},z)\to 0,\ \text{as}\ n\to\infty,$
from triangle inequality
$d(x_{\mu+1},x_{g(n)+1})-d(x_{\mu+1},z)$ $\leq d(x_{g(n)+1},z),$
$d(x_{\mu+1},z)-d(x_{\mu+1},x_{g(n)+1})$ $\leq d(x_{g(n)+1},z).$
From the first equality in (3.4), $d(x_{\mu+1},z)\not=0$ holds. By (3.4) we get
$\frac{1}{2}d(x_{\mu+1},z)\leq\frac{1}{2}d(x_{\mu},z).$
Then, $d(x_{\mu},z)\not=0$ holds. So by (CM) we obtain
$d(x_{\mu+1},z)=d(Tx_{\mu},Tz)<\frac{1}{2}\{d(x_{\mu},x_{\mu+1})+d(z,Tz)\}=\frac{1}{2}d(x_{\mu},x_{\mu+1}),$
which is a contradiction with (3.4). Therefore (i) holds. ∎
4. In the case of Chatterjea type contraction mapping
We similarly show a fixed point theorem and weakest condition in the CJM version for contraction mappings of Chatterjea type.
Theorem 4.1.
Let $(X,d)$ be a complete metric space and suppose that the mapping $T:X\to X$ satisfies the following.
  1. (i)
    For every $\varepsilon>0$, there exists $\delta>0$ such that $\frac{1}{2}\{d(x,Ty)+d(y,Tx)\}<\varepsilon+\delta$ implies $d(Tx,Ty)\leq\varepsilon$.
  2. (ii)
    $x\not=y$ implies $d(Tx,Ty)<\frac{1}{2}\{d(x,Ty)+d(y,Tx)\}$.
Then, $T$ has a unique fixed point.
Proof.
Fix $x_{0}\in X$. Set $x_{n}=T^{n}x_{0}$ for all $n\in\mathbb{N}$. If there exists $N\in\mathbb{N}$ such that $x_{N}=x_{N+1}$, $x_{N}$ is a fixed point of $T$. Therefore, assume that $x_{n}\not=x_{n+1}$ for any $n\in\mathbb{N}$. Then, we have
$d(x_{n},x_{n+1})$ $=d(Tx_{n-1},Tx_{n})<\frac{1}{2}\{d(x_{n-1},Tx_{n})+d(x_{n},Tx_{n-1})\}$
$=\frac{1}{2}\{d(x_{n-1},x_{n+1})+d(x_{n},x_{n})\}\leq\frac{1}{2}\{d(x_{n-1},x_{n})+d(x_{n},x_{n+1})\}.$
So,
$\frac{1}{2}d(x_{n},x_{n+1})<\frac{1}{2}d(x_{n-1},x_{n}).$
Therefore, $\{d(x_{n},x_{n+1})\}$ is a strictly decreasing sequence.
Thus, there exists $\varepsilon_{0}\geq 0$ such that $d(x_{n},x_{n+1})\to\varepsilon_{0}$ as $n\to\infty$ and $d(x_{n},x_{n+1})>\varepsilon_{0}$ for any $n$.
Assume that $\varepsilon_{0}>0$. Then there exist $\delta_{0}>0$ and $k\in\mathbb{N}$ such that
$\frac{1}{2}\{d(x_{k},Tx_{k+1})+d(x_{k+1},Tx_{k})\}\leq\frac{1}{2}\{d(x_{k},x_{k+1})+d(x_{k+1},Tx_{k+1})\}<\varepsilon_{0}+\delta_{0}.$
By assumption (i) we have $d(Tx_{k},Tx_{k+1})\leq\varepsilon_{0}$, which is a contradiction. Therefore, $\varepsilon_{0}=0$ and
(4.1) $\lim_{n\to\infty}d(x_{n},x_{n+1})=0.$
Next, we show that $\{x_{n}\}$ is a Cauchy sequence. Let $\varepsilon>0$ be arbitrary and let $\delta$ be chosen such that $\delta\leq\varepsilon$. By (4.1) there exists some $k\in\mathbb{N}$ such that $d(x_{n-1},x_{n})<\delta$ for each $n>k$. Fix $n>k$. It suffices to show that
$d(x_{n},x_{n+p})\leq\varepsilon,\ \text{for all}\ p\in\mathbb{N}.$
We show by induction. The case of $p=1$ is obvious. We show the case of $p$ assuming the case of $p-1$. In this case we have
$\frac{1}{2}\{d(x_{n-1},Tx_{n+p-1})+d(x_{n+p-1},Tx_{n-1})\}=\frac{1}{2}\{d(x_{n-1},x_{n+p})+d(x_{n+p-1},x_{n})\}$
$\leq\frac{1}{2}\{d(x_{n-1},x_{n})+d(x_{n},x_{n+p-1})+d(x_{n+p-1},x_{n+p})+d(x_{n+p-1},x_{n})\}$
$<\delta+\varepsilon.$
Therefore,
$d(x_{n},x_{n+p})=d(Tx_{n-1},Tx_{n+p-1})\leq\varepsilon.$
Then since $\{x_{n}\}$ is a Cauchy sequence, there exists $u\in X$ such that $x_{n}\to u$ as $n\to\infty$. From
$d(u,Tu)$ $\leq d(u,x_{n})+d(x_{n},Tx_{n})+d(Tx_{n},Tu)$
$<d(u,x_{n})+d(x_{n},Tx_{n})+\frac{1}{2}\{d(x_{n},Tu)+d(u,Tx_{n})\}$
$\leq d(u,x_{n})+d(x_{n},Tx_{n})+\frac{1}{2}d(x_{n},Tu)+\frac{1}{2}d(u,x_{n})+\frac{1}{2}d(x_{n},Tx_{n})$
$\leq d(u,x_{n})+d(x_{n},Tx_{n})+\frac{1}{2}\{d(x_{n},u)+d(u,Tu)\}+\frac{1}{2}d(u,x_{n})+\frac{1}{2}d(x_{n},Tx_{n}),$
we have
$\frac{1}{2}d(u,Tu)<2d(u,x_{n})+\frac{3}{2}d(x_{n},Tx_{n})\to 0\ (n\to\infty).$
Then we get $u=Tu$.
For uniqueness, let $u\not=v,\ Tu=u$ and $Tv=v$. Then we have
$d(u,v)$ $=d(Tu,Tv)<\frac{1}{2}\{d(u,Tv)+d(v,Tu)\}$
$\leq\frac{1}{2}\{d(u,v)+d(v,Tv)\}+\frac{1}{2}\{d(v,u)+d(u,Tu)\}=d(u,v),$
which is a contradiction. ∎
Theorem 4.2.
Let $(X,d)$ be a complete metric space and suppose that the mapping $T:X\to X$ satisfies the following.
(CM2) $x\not=y\ \text{implies}\ d(Tx,Ty)<\frac{1}{2}\{d(x,Ty)+d(y,Tx)\}.$
Then, the following are equivalences.
  1. (i)
    For any $x\in X$ and $\varepsilon>0$, there exists $\delta>0$ such that
    $\frac{1}{2}\{d(T^{i}x,T^{j+1}x)+d(T^{j}x,T^{i+1}x)\}<\varepsilon+\delta$ implies $d(T^{i+1}x,T^{j+1}x)\leq\varepsilon$ for all $i,j\in\mathbb{N}\cup\{0\}$.
  2. (ii)
    $T$ has a unique fixed point $z\in X$ and $\{T^{n}x\}$ converges to $z$ for any $x\in X$.
Proof.
((i)$\Rightarrow$(ii)) It can be shown as in the proof of Theorem 4.1.
((ii)$\Rightarrow$(i)) Let $z\in X$ be a fixed point of $T$. Arguing by contradiction, we assume that (i) does not hold, that is, there exist $x\in X$ and $\varepsilon>0$ such that for all $n\in\mathbb{N}$ there exists $\{f(n)\},\{g(n)\}\subset\mathbb{N}\cup\{0\}$ such that
$f(n)<g(n),$
(4.2) $\frac{1}{2}\{d(x_{f(n)},x_{g(n)+1})+d(x_{g(n)},x_{f(n)+1})\}<\varepsilon+\frac{1}{n}\quad\text{and}$
(4.3) $\varepsilon<d(x_{f(n)+1},x_{g(n)+1}),$
where $x_{f(n)}=T^{f(n)}x$. Without loss of generality, we may assume that either
  • $\lim_{n\to\infty}f(n)=\infty$  or
  • There exists $\mu\in\mathbb{N}\cup\{0\}$ such that $f(n)=\mu$ for all $n\in\mathbb{N}$.
Note that if $\{f(n)\}$ oscillates, the second case is attributed. Similarly, we may assume that either
  • $\lim_{n\to\infty}g(n)=\infty$  or
  • There exists $\nu\in\mathbb{N}\cup\{0\}$ such that $g(n)=\nu$ for all $n\in\mathbb{N}$.
By (4.3) we have $x_{f(n)}\not=x_{g(n)}$. So, by (CM2) we obtain
$\varepsilon$ $<d(x_{f(n)+1},x_{g(n)+1})<\frac{1}{2}\{d(x_{f(n)},x_{g(n)+1})+d(x_{g(n)},x_{f(n)+1})\}<\varepsilon+\frac{1}{n}.$
Therefore, we get
(4.4) $0<\varepsilon=\lim_{n\to\infty}d(x_{f(n)+1},x_{g(n)+1})=\frac{1}{2}\lim_{n\to\infty}\{d(x_{f(n)},x_{g(n)+1})+d(x_{g(n)},x_{f(n)+1})\}.$
Since $\{x_{n}\}$ converges to $z$, $f(n)\not\to\infty$ or $g(n)\not\to\infty$ holds. Suppose that $f(n)\not\to\infty$ holds, that is, $f(n)=\mu$ for all $n\in\mathbb{N}$ is satisfied.
First, if $g(n)=\mu$ for all $n\in\mathbb{N}$, we have
$d(x_{\mu+1},x_{\nu+1})=\varepsilon,$
which is a contradiction. On the other hand, if $\lim_{n\to\infty}g(n)=\infty$, then $x_{g(n)}\to z$, so from (4.4) we have
$d(x_{\mu+1},z)=\varepsilon$ $=\frac{1}{2}\lim_{n\to\infty}\{d(x_{f(n)},x_{g(n)+1})+d(x_{g(n)},x_{f(n)+1})\}$
(4.5) $=\frac{1}{2}\{d(x_{\mu},z)+d(z,x_{\mu+1})\}.$
We can show the third equality in (4) since we obtain
|$|\{d(x_{\mu},z)+d(z,x_{\mu+1})\}-\{d(x_{\mu},x_{g(n)+1})+d(x_{g(n)},x_{\mu+1})\}|$
$\leq|d(x_{g(n)},z)|+|d(x_{g(n)+1},z)|\to 0,\ \text{as}\ n\to\infty,$
from triangle inequality
$d(x_{\mu+1},z)-d(x_{\mu+1},x_{g(k)})$ $\leq d(x_{g(k)},z),$
$d(x_{\mu},z)-d(x_{\mu},x_{g(k)+1})$ $\leq d(x_{g(k)+1},z).$
By (4), $d(x_{\mu},z)\not=0$ holds. So by (CM2) we obtain
$d(x_{\mu+1},z)=d(Tx_{\mu},Tz)<\frac{1}{2}\{d(x_{\mu},Tz)+d(z,Tx_{\mu})\}=\frac{1}{2}\{d(x_{\mu},z)+d(z,x_{\mu+1})\},$
which is a contradiction with (4). Therefore (i) holds. ∎
Declarations
Conflicts of interests
The authors declare that there is no conflict of interest regarding the publication of this paper.
Data Availability Statements
Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.